OpenStudy (anonymous):

If the functions f(x)=ax^3-bx+1 and g(x)=x^2-2ax-b for x=1 the vertical length beetween the functions is 3 and Cf and Cg intersect with the line ε=2x+1 find a,b

OpenStudy (anonymous):

Plugging in x = 1 into both f and g we get: \[f(1) = a-b+1, g(1) = 1-2a-b\] now it doesnt say which function is greater, so im going to assume f(x) > g(x) and hope for the best. Since the vertical length between them is 3 at x = 1, we subtract them and set it equal to three: \[f(1)-g(1) = 3 \Rightarrow (a-b+1)-(1-2a-b) = 3 \Rightarrow 3a = 3 \Rightarrow a = 1\]

myininaya (myininaya):

f(1)=a-b+1 g(1)=1-2a-b \[d=\sqrt{(f(1)-g(1))^2}=3\] \[f(1)-g(1)=3\] \[(a-b+1)-(1-2a-b)=3\]

myininaya (myininaya):

lol gj joe

OpenStudy (anonymous):

lol, i was on the phone and had a bajillion mistakes in it >.<

OpenStudy (anonymous):

looks good to me. what the monkey is Cf and Cg?

myininaya (myininaya):

im not sure

OpenStudy (anonymous):

a random constant times the function was my guess, but i havent started working it out yet

OpenStudy (anonymous):

ima do it on paper before i type junk >.>

OpenStudy (anonymous):

can't wait to see this...

myininaya (myininaya):

so i guess we set Cf=2x+1 and set Cg=2x+1

myininaya (myininaya):

i wonder if we should assume both constants are the same i think not

OpenStudy (anonymous):

Cf and Cg for the graph of f(x) and g(x)

myininaya (myininaya):

c(x^3-bx+1)=2x+1 and we have c'(x^2-2b-b)=2x+1

OpenStudy (anonymous):

Thank you very much!! this is a tough one for me!

myininaya (myininaya):

hey jason what does C mean is that a constant times the function?

OpenStudy (anonymous):

yeah, after playing with it a little, i dont think its a constant anymore lol

OpenStudy (anonymous):

i thing it might be craft, i am not sure,that is how we symbolize the graph of a function in greece!!!

myininaya (myininaya):

\[x^3-bx-2x=0\] \[x^3-x(b+2)=0\] \[x(x^2-(b+2))=0\] \[x=0, x^2=b+2\] \[x=0, x= \pm \sqrt{b+2}\] \[x^2-2x-3b-1=0\] \[x=0=> -3b-1=0\] \[b=\frac{1}{-3}\]

myininaya (myininaya):

we have two equations above since both f and g intersect the line e=2x+1

myininaya (myininaya):

so i did f-(2x+1)=0 and g-(2x+1)=0

OpenStudy (anonymous):

are you making this up?

OpenStudy (anonymous):

ah, very nice

OpenStudy (anonymous):

dont we all "make it up"? :P gotta go with the flow

myininaya (myininaya):

solved f-(2x+1)=0 this for x and then plugged x into g-(2x+1)=0 to solve for b (and we already knew a=1 from joe)

myininaya (myininaya):

also i plug the pretty x into g-(2x+1)=0 not the scary x

OpenStudy (anonymous):

Geniusly done!

OpenStudy (anonymous):

its definitely correct, i tested those values of a and b to see if at x =1 they were 3 units apart, its good.

OpenStudy (anonymous):

\[\color{blue}{\text{myininaya is a hero}}\]

myininaya (myininaya):

let me know if you have any questions jason?

OpenStudy (anonymous):

no understood everything,but just to let you know this exersise is just one of the simple ones (let's say) for our final exams in next years exams in may,so this is just the a start,all this to a 17 years old kids,Anyway, THANK YOU VERY MUCH FOR YOUR HELP

myininaya (myininaya):

this looks like advanced algebra lol the college algebra i seen in schools don't ask questions like this

myininaya (myininaya):

oh you are 17 lol

myininaya (myininaya):

i wish i was challenge more in high school i hate my high school

OpenStudy (anonymous):

WELL here in greece the last highschool year is really tough!we will be solving exersises with function aprox. 3 books, and make over 30 of practice exams

myininaya (myininaya):

hated and hate*

OpenStudy (anonymous):

seems intense o.O its worth it though

OpenStudy (anonymous):

yeah I also wish my high school was more challenging. would have prepared me a lot more for college and beyond.